OpenStudy (anonymous):

Did I do this right? SA= 2(3 ½ + 11 ½ ) + 2(11 ½ + 11 ¾ ) + 2(3 ½ + 11 ¾ ) SA = 2(30) + 2(46.5) + 2(30.5) SA= 60 + 93 + 61 SA= 214

OpenStudy (anonymous):

It looks like you already multiplied everything by two in the second step and put it in parenthesis. So, the additional 2 outside would not be needed at this point.

OpenStudy (anonymous):

SA = 2(15) + 2(23.25) + 2(15.25) should be what you have in step 2.

OpenStudy (anonymous):



OpenStudy (anonymous):

Basically step 2 should be SA = 2(15) + 2(23.25) + 2(15.25).

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

So all i have to do is do SA = 2(15) + 2(23.25) + 2(15.25) and get my answer?

OpenStudy (anonymous):

It looked like you accidentally multiplied by 2 going to that step and didn't realize it. Otherwise it would be (30) + (46.5) + (30.5). But that would require two steps.

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

ok let me see what i get

OpenStudy (anonymous):

ok 107?

OpenStudy (anonymous):

That is correct. Also there is another way you could have done it which would give the same answer. Since everything is multiplied by 2, you can separate it and add everything else together.

OpenStudy (anonymous):

yea i think i would have liked it better that way

OpenStudy (anonymous):

SA= 2(3 ½ + 11 ½ ) + 2(11 ½ + 11 ¾ ) + 2(3 ½ + 11 ¾ ) SA = 2(3 ½ + 11 ½ + 11 ½ + 11 ¾ + 3 ½ + 11 ¾ ) SA= 2(53 ½) SA= 107

OpenStudy (anonymous):

i have another question can you just explain it to me what it means?

OpenStudy (anonymous):

4. What is the surface area of the box if it is scaled up by a factor of 10? its with the same question i asked before

OpenStudy (anonymous):

I dont know what it wants me to do

OpenStudy (anonymous):

Scaling something by a factor essentially means multiplying everything in the equation by that factor.

OpenStudy (anonymous):

Since your factor is 10, you would simply multiply the Surface Area you already have by that factor (10).

OpenStudy (anonymous):

so 107/10?

OpenStudy (anonymous):

107 * 10

OpenStudy (anonymous):

Out of curiosity, what was the original question that brought you to your work that you originally asked about?

OpenStudy (anonymous):

1,070 right?

OpenStudy (anonymous):

1. List the dimensions of your box. Be sure to include the units (in, cm, ft., etc.). 2. Describe the shape of the cross section when the box is cut parallel to the base. 3. What is the surface area of the box? 4. What is the surface area of the box if it is scaled up by a factor of 10? 5. What is the volume of the box? 6. What is the volume of the box if it is scaled down by a factor of ½?

OpenStudy (anonymous):

OK, we need to back up to your original question as you set up the formula wrong.

OpenStudy (anonymous):

ok ok #work lol

OpenStudy (anonymous):

hehe, it happens. Did you have to pick an item to measure or was dimensions already give to you?

OpenStudy (anonymous):

i had to pick it

OpenStudy (anonymous):

i just chose a binder

OpenStudy (anonymous):

The formula for SA is 2 ( lw + lh + hw) or 2(lw) + 2(lh) + 2(hw) where the letters besides each other mean to multiply, not add.

OpenStudy (anonymous):

Sorry, I didn't think about your equation when you asked, was just making sure the math was correct.

OpenStudy (anonymous):

So, you would be multiply fractions by using those measurements.

OpenStudy (anonymous):

Would you like to try something different or keep with the binder?

OpenStudy (anonymous):

um i already have the measurements so keep it lol

OpenStudy (anonymous):

Ok, so now you would have SA= 2(3 ½ * 11 ½ ) + 2(11 ½ * 11 ¾ ) + 2(3 ½ * 11 ¾ )

OpenStudy (anonymous):

135.125?

OpenStudy (anonymous):

SA= 2(3 ½ * 11 ½ ) + 2(11 ½ * 11 ¾ ) + 2(3 ½ * 11 ¾ ) SA= 2(40.25 ) + 2(135.125 ) + 2(41.125 )

OpenStudy (anonymous):

sorry, this place is annoying with fractions. .125 would be 1/8